A Candy store sells red jelly beans and blue jelly beans Caden buys 8 ounces of red and 4 of blue jelly beans for a total of $20. Tovah buys 6 ounces of red and 9 of blur jelly beans for a total of $27. What is the cost per ounce for red and blue jelly beans?
Answer by josgarithmetic(39615)

Price of red jelly beans, r.
Price of blue jelly beans, b.
Price unit is dollars per ounce.

Caden's cost: 8r+4b=20
Tovah's cost: 6r+9b=27

Each of the equations can and should be simplified. Divide Caden's equation by 4, and divide Tovah's equation by 3.

Caden: 2r+b=5
Tovah: 2r+3b=9

You could use substitution to solve this system, but elimination will be easier, and maybe faster. SUBTRACT Caden equation from Tovah equation.
(2r+3b)-(2r+b)=9-5
b=4
-
Finding r price of red jelly beans,
2r+b=5
2r=5-b
r=(5-b)/2
r=(5-4)/2
r=1/2 dollars per ounce, Red Jelly Beans.
-
-
--------------------------------
Blue Jelly Beans, $4/ounce
Red Jelly Beans, $0.50/ounce
--------------------------------